# On sporadic symmetry breaking operators for principal series representations of the de Sitter and Lorentz groups

### v1

@@ 4.3 Phase 3: Proving $f_{-0}=f_{+0}$
$f_{-j}\equiv 0\text{ for all }j=0,\ldots,N,\quad\text{and}\quad f_{j}\equiv 0\text{ for all }j=0,\ldots,m-a-1.$
Thus, a solution has the form $(0,\ldots,0,f_{m-a},\dots,f_{N})$. By Proposition 4.11, this solution must satisfy (4.15). Thus, by (4.2), the terms of degree greater than $a-m+j$ appearing in the expression (4.15) of $f_{j}$ must vanish. In particular, the term $t^{a-m+j+2}$ must be zero. A direct computation by using the definition of the Gegenbauer polynomial shows that this term is equal to
$\enspace\frac{\Gamma(N+j+1)}{\Gamma(N+1)}\frac{(-1)^{m-N-1}2^{j}(2i)^{a-m+2}}{(a-m+j+2)!(m-N-1)!}\frac{\Gamma\left(\lambda+a+j\right)}{\Gamma\left(\lambda+\left[\frac{a+m-1}{2}\right]\right)}$
×$\times(N+j+1)_{N-j}\sum_{r=0}^{N-j}\frac{(j-N)_{r}(\lambda+a-N-1)_{r}}{(-2N)_{r}r!}.$
@@
First, we note that for any $j=0,\ldots,N$, the expression (4.43) has degree less than or equal to $a-m+j$. In fact, fix $q$ and think about the degree of the term $D_{j,r}\widetilde{C}_{a-m+j+(q-2N+r)}^{\lambda+N-1-r}$ for a given $r$. If $q-2N+r\leq 0$, then $a-m+j+2(q-2N+r)\leq a-m+j$, so the term has degree $\leq a-m+j$. On the contrary, if $q-2N+r>0$, then $D_{j,r}=0$ as $\frac{\Gamma(q-2N+r)}{\Gamma(q-2N)}=0$. Thus, all terms in the sum have degree $\leq a-m+j$. In particular, this expression for $f_{j}$ satisfies (4.2).
@@
In fact, set $\ell:=a+m-2N+j+2r$ and $b:=-\lambda-N+1+r=a+q-2N+r$. Since $q\geq\max(0,N-a+m)$ we have $\ell,b\in{\mathbb{N}}$. Now, if $2b-\ell=a-m+2q-j-2N\geq 0$, then (4.45) holds thanks to (8.10). On the other hand, if $02b-\ell=a-m+2q-j-2N<$ 0 the right-hand side of (4.45) vanishes. But so does the left-hand side as
@@
By a change of indexes, the sum part can be written as
@@
appearing in (4.43) is a well-defined non-zero constant for any $a\geq m-N$, any $j=0,\ldots,N$ and any $\lambda=N+1-a-q\in\Lambda(N,a)$. In fact, if $j-2N+2q=0$, (4.46) equals to $1$. If $j-2N+2q\geq 1$, then (4.46) amounts to
@@
5 The case $m<-N$
In this section we recall a duality theorem from [36] that shows the existence of an isomorphism between the F-systems $\text{{Sol}}(\mathfrak{n}_{+};\sigma^{2N+1}_{\lambda},\tau_{m,\nu})$ and $\text{{Sol}}(\mathfrak{n}_{+};\sigma^{2N+1}_{\lambda},\tau_{-m,\nu})$ when $|m|\geq N$. This allows us to deduce the case $m<-N$ from the case $m>N$ in Theorems 1.3 and 1.4.
@@ 5 The case $m<-N$
is a well-defined isomorphisms.
With the result above it is clear that we can deduce the case $m<-N$ from the case $m>N$ in Theorems 1.3 and 1.4. We write the detailed proofs below.
Proof of Theorem 1.3 for $m<-N$.
@@
Proof of Theorem 1.4 for $m<-N$.
From Proposition 5.1 and the definition of the map $\Phi^{+}_{\text{Sol}}$, the expression of the operator $\mathbb{D}_{\lambda,\nu}^{N,m}$ can be obtained by changing the coordinate $u_{s}$ by the coordinate $(-1)^{s}u_{2N-s}$ (for $s=0,\ldots,2N$), and by switching $i$ by $i$ (or, equivalently, $z$ by $\overline{z}$) in the expression of the operator $\mathbb{D}_{\lambda,\nu}^{N,-m}$. By doing the change of index given by $d=2N-d$, we obtain precisely (1.6). ∎
@@ 6 Localness Theorem
In this section we prove the localness theorem (Theorem 1.6). The proof is based in the work [27, 28] and uses several known results about distributions and symmetry breaking operators.
First, recall from [28, Fact 5.1] that there is a natural bijection
@@
$\left(\pi\begin{pmatrix}1&&\\ &g&\\ &&1\end{pmatrix}F_{1}\right)(a)$ $=(\sigma^{2N+1})^{\vee}(g)F_{1}(g^{-1}a),\quad\text{for }g\in SO(3),$
$\left(\pi(e^{tH_{0}})F_{1}\right)(a)$ $=e^{(\lambda-3)t}F_{1}(e^{-t}a),\quad\text{for }t\in{\mathbb{R}},$
$\left(\pi(n_{+}(c))F_{1}\right)(a)$ $=F_{1}(a-c),\quad\text{for }c\in{\mathbb{R}}^{3}.$
@@
One can describe the $P$-action on the second projection space as well. However, in this case, since $U_{2}=N_{-}P/P$ is not $N_{+}$-invariant, the action of $N_{+}$ shall be substituted by its infinitesimal action on the Lie algebra $\mathfrak{n}_{+}$. We omit a detailed description of this action, but it can be verified in [28, Lem. 5.13].
@@
[…]
@@
This result can be proved following a similar argument to that of [28, Thm. 3.6].
Recall from [36, Lem. 4.1] that the representation $V^{2N+1}$ can be decomposed as a $SO(2)$-module as follows:
@@
In this section we prove Theorem 7.3, that assures that any non-zero SBO in (1.1) for $|m|>N$ is sporadic in the sense of Kobayashi–Speh (see Definition 7.1). We prove this result by linking the space of SBOs for the connected pair $(SO_{0}(4,1),SO_{0}(3,1))$ to that of the non-connected pair $(O(4,1),O(3,1))$, and then use the known results about sporadicity for the latter. Let us start by fixing some notation, which is slightly different from other sections of the paper.
We denote by $(G,G^{\prime})$ the non-connected pair $(O(4,1),O(3,1))$ whereas $(G_{0},G_{0}^{\prime})$ will denote the identity component of $(G,G^{\prime})$, namely:
[…]
Following the geometric realization given in Section 2, it is clear that $G$ acts transitively on $\Xi/{\mathbb{R}}^{\times}$, and that $P:=\text{{stab}}_{G}(1:0:0:0:1)$ is a parabolic subgroup of $G$ satisfying
$G/P\simeq G_{0}/P_{0}\simeq\Xi/{\mathbb{R}}^{\times}\simeq S^{3},$
where $P_{0}$ denotes the identity component of $P$. The same notation by using zero as a subscript will be used for other subgroups as well. The Langlands decomposition of $P$ is given as follows:
$P=MAN_{+},$
$M=\left\{\begin{pmatrix}\varepsilon&&\\ &g&\\ &&\varepsilon\end{pmatrix}:g\in O(3),\varepsilon=\pm 1\right\}\simeq O(3)\times O(1),$
$A=\exp({\mathbb{R}}H_{0})\simeq SO_{0}(1,1),$
$N_{+}=\exp(\mathfrak{n}_{+}({\mathbb{R}}))\simeq{\mathbb{R}}^{3},$
and the Langlands decomposition of $P_{0}$ given in Section 2 is clearly $P_{0}=M_{0}AN_{+}$.
Similarly as before, $G^{\prime}$ can be realized as $G^{\prime}\simeq\text{{stab}}_{G}{}^{t}(0,0,0,1,0)$, and $P^{\prime}=P\cap G^{\prime}$ defines a parabolic subgroup of $G^{\prime}$ satisfying
$G^{\prime}/P^{\prime}\simeq G_{0}^{\prime}/P_{0}^{\prime}\simeq(\Xi\cap\{x_{3}=0\})/{\mathbb{R}}^{\times}\simeq S^{2}.$
The Langlands decomposition of $P^{\prime}$ is given as follows:
$P^{\prime}=M^{\prime}AN_{+}^{\prime},$
$M^{\prime}=\left\{\begin{pmatrix}\varepsilon&&&\\ &g&&\\ &&1\\ &&&\varepsilon\end{pmatrix}:g^{\prime}\in O(2),\varepsilon=\pm 1\right\}\simeq O(2)\times O(1),$
$N_{+}^{\prime}=\exp(\mathfrak{n}_{+}({\mathbb{R}}))\simeq{\mathbb{R}}^{2}.$
And clearly, the Langlands decomposition of $P^{\prime}_{0}$ is given by $P^{\prime}_{0}=M_{0}^{\prime}AN_{+}^{\prime}$.
Now, given $(\sigma,V)\in\widehat{O(3)}$, $\alpha\in\widehat{O(1)}\simeq{\mathbb{Z}}/2{\mathbb{Z}}$ and $\lambda\in{\mathbb{C}}$, we denote by $I(V,\lambda)_{\alpha}$ the following principal series representation of $G$
$I(V,\lambda)_{\alpha}:=\text{{Ind}}_{P}^{G}(\sigma\boxtimes(-1)^{\alpha}\boxtimes{\mathbb{C}}_{\lambda}),$
where $(-1)^{\alpha}$ is the character of $O(1)\subset M$ given by the trivial representation for $\alpha=0$, and the sign representation for $\alpha=1$.
Similarly, we define the principal series representation $J(W,\nu)_{\beta}$ of $G^{\prime}$ for $(\tau,W)\in\widehat{O(2)}$, $\beta\in\widehat{O(1)}\simeq{\mathbb{Z}}/2{\mathbb{Z}}$ and $\nu\in{\mathbb{C}}$ as follows
$J(W,\nu)_{\beta}:=\text{{Ind}}_{P^{\prime}}^{G^{\prime}}(\tau\boxtimes(-1)^{\beta}\boxtimes{\mathbb{C}}_{\nu}).$
The space of SBOs for $(G,G^{\prime})$ between these two principal series representations is given as follows:
$\text{{Hom}}_{G^{\prime}}(I(V,\lambda)_{\alpha},J(W,\nu)_{\beta}).$ (7.1)
For the connected pair, we define the following induced representations
$I_{0}(V_{0},\lambda):=\text{{Ind}}_{P_{0}}^{G_{0}}(\sigma_{0}\boxtimes{\mathbb{C}}_{\lambda}),$
$J_{0}(W_{0},\nu):=\text{{Ind}}_{P_{0}^{\prime}}^{G_{0}^{\prime}}(\tau_{0}\boxtimes{\mathbb{C}}_{\nu}),$
where $(\sigma_{0},V_{0})$ and $(\tau_{0},W_{0})$ are finite-dimensional representations of $SO(3)$ and $SO(2)$, respectively. The space of SBOs for $(G_{0},G_{0}^{\prime})$ between the two induced representations above is given as follows:
$\text{{Hom}}_{G_{0}^{\prime}}(I_{0}(V_{0},\lambda),J_{0}(W_{0},\nu)).$ (7.2)
For the pair $(V_{0},W_{0})=(V^{2N+1},{\mathbb{C}}_{m})$ defined in Section 2, it is clear that
$I_{0}(V^{2N+1},\lambda)=C^{\infty}(S^{3},\mathcal{V}_{\lambda}^{2N+1}),$
$J_{0}({\mathbb{C}}_{m},\nu)=C^{\infty}(S^{2},\mathcal{L}_{m,\nu}).$
Therefore, for this pair, (7.2) is nothing but the space (1.1).
Now, we define sporadicity.
@@
Given $V_{0}\simeq V^{2N+1}\in\widehat{SO(3)}$ and $W_{0}\simeq{\mathbb{C}}_{m}\in\widehat{SO(2)}$, let $\mathcal{P}\equiv\mathcal{P}_{V_{0},W_{0}}\subset{\mathbb{C}}^{2}$ be the set of parameters $(\lambda,\nu)$ such that the space of SBOs (1.1) is non-zero. For a fixed $(N,m)$ and $(\lambda,\nu)\in\mathcal{P}$, we say that an SBO ${\mathbb{T}}_{\lambda,\nu}^{N,m}$ is sporadic with respect to the pair $(G_{0},G_{0}^{\prime})$ if $(\lambda,\nu)$ is an isolated point of $\mathcal{P}$. In other words, if there is an open neighbourhood $U$ of $(\lambda,\nu)$ on ${\mathbb{C}}^{2}$ such that $U\cap\mathcal{P}=\{(\lambda,\nu)\}$.
Although Definition 7.1 above is stated for the pair $(G_{0},G_{0}^{\prime})$, the notion of sporadicity can be formulated also for the pair $(G,G^{\prime})$ in the obvious way.
@@
    Note that the notion of sporadicity given in [28] by Kobayashi–Speh is slightly different to the one defined above. However, it is straightforward to check that their definition implies ours.
@@
In order to prove Theorem 7.3 above, we make use of two preliminary results. One gives a characterization of the sporadic SBOs for the non-connected pair (Proposition 7.4) and the other relates the space of SBOs for the connected pair to that for the non-connected pair (Proposition 7.5).
Proposition 7.4 ([28, Thm. 3.9]).
Let $\lambda,\nu\in{\mathbb{C}}$, $\alpha,\beta\in{\mathbb{Z}}/2{\mathbb{Z}}$ and $V\in\widehat{O(3)},W\in\widehat{O(2)}$ and suppose that the space of SBOs (7.1) for the pair $(O(4,1),O(3,1))$ is non-zero. Then, the following assertions are equivalent:
  1. (i)
    $[V:W]:=\dim_{\mathbb{C}}\text{{Hom}}_{O(2)}(V\big{\rvert}_{O(2)},W)=0$.
  2. (ii)
    Any SBO in (7.1) is sporadic.
Now, for $N,m\in{\mathbb{N}}$, consider $(\sigma,V)\in\widehat{O(3)}$ and $(\tau,W)\in\widehat{O(2)}$ given by the natural representations on the spaces of harmonic polynomials $V=\mathcal{H}^{N}({\mathbb{C}}^{3})$ and $W=\mathcal{H}^{m}({\mathbb{C}}^{2})$, respectively. Then, we note that
$\mathcal{H}^{N}({\mathbb{C}}^{3})\big{\rvert}_{SO(3)}\simeq V^{2N+1},$
$\mathcal{H}^{m}({\mathbb{C}}^{2})\big{\rvert}_{SO(2)}\simeq{\mathbb{C}}_{m}\oplus{\mathbb{C}}_{-m},\quad(m>0).$
Therefore, we have natural isomorphisms
$I(\mathcal{H}^{N}({\mathbb{C}}^{3}),\lambda)_{\alpha}\big{\rvert}_{SO_{0}(4,1)}\simeq I_{0}(V^{2N+1},\lambda),$ (7.3)
$J(\mathcal{H}^{m}({\mathbb{C}}^{2}),\nu)_{\beta}\big{\rvert}_{SO_{0}(3,1)}\simeq J_{0}({\mathbb{C}}_{m},\nu)\oplus J_{0}({\mathbb{C}}_{-m},\nu),\quad(m>0).$
Moreover, we obtain the following.
Proposition 7.5.
Suppose $\lambda,\nu\in{\mathbb{C}}$, $N\in{\mathbb{N}}$ and $m\in{\mathbb{N}}_{+}$. Then,
  1. (1)
    There is a natural isomorphism
    $\text{{Hom}}_{SO_{0}(3,1)}(I_{0}(\mathcal{H}^{N}({\mathbb{C}}^{3}),\lambda),J_{0}$ $(\mathcal{H}^{m}({\mathbb{C}}^{2}),\nu))$
    $\simeq\text{{Hom}}_{SO_{0}(3,1)}(I_{0}(V^{2N+1},\lambda),J_{0}({\mathbb{C}}_{m},\nu))$
    $\oplus\text{{Hom}}_{SO_{0}(3,1)}(I_{0}(V^{2N+1},\lambda),J_{0}({\mathbb{C}}_{-m},\nu)).$
  2. (2)
    Further, suppose $\nu-\lambda\in{\mathbb{N}}$ and $m>N$. Fix $\alpha\in{\mathbb{Z}}/2{\mathbb{Z}}$ and set $\beta:=\alpha+\nu-\lambda\mod 2$. Then, we have
    $\text{{Hom}}_{SO_{0}(3,1)}(I_{0}(\mathcal{H}^{N}({\mathbb{C}}^{3}),\lambda)$ $,J_{0}(\mathcal{H}^{m}({\mathbb{C}}^{2}),\nu))$ (7.4)
    $\simeq\text{{Hom}}_{O(3,1)}(I(\mathcal{H}^{N}({\mathbb{C}}^{3}),\lambda)_{\alpha},J(\mathcal{H}^{m}({\mathbb{C}}^{2}),\nu)_{\beta})$
    $\oplus\text{{Hom}}_{O(3,1)}(I(\mathcal{H}^{N}({\mathbb{C}}^{3}),\lambda)_{\alpha},J(\mathcal{H}^{m}({\mathbb{C}}^{2})\otimes\det,\nu)_{\beta}).$

### v2

@@ 4.3 Phase 3: Proving $f_{-0}=f_{+0}$
$f_{-j}\equiv 0\text{ for all }j=0,\ldots,N,\qquad\text{and}\qquad f_{j}\equiv 0\text{ for all }j=0,\ldots,m-a-1.$
Thus, a solution has the form $(0,\ldots,0,f_{m-a},\dots,f_{N})$. By Proposition 4.11, this solution must satisfy (4.15). Thus, by (4.2), the terms of degree greater than am+ja-m+j appearing in the expression (4.15) of $f_{j}$ must vanish. In particular, the term $t^{a-m+j+2}$ must be zero. A direct computation by using the definition of the Gegenbauer polynomial shows that this term is equal to
$\kern 4.62497pt\frac{\Gamma(N+j+1)}{\Gamma(N+1)}\frac{(-1)^{m-N-1}2^{j}(2i)^{a-m+2}}{(a-m+j+2)!(m-N-1)!}\frac{\Gamma\left(\lambda+a+j\right)}{\Gamma\left(\lambda+\left[\frac{a+m-1}{2}\right]\right)}$
$\times(N+j+1)_{N-j}\sum_{r=0}^{N-j}\frac{(j-N)_{r}(\lambda+a-N-1)_{r}}{(-2N)_{r}r!}.$
@@
First, we note that for any $j=0,\ldots,N$, the expression (4.43) has degree less than or equal to am+ja-m+j. In fact, fix $q$ and think about the degree of the term $D_{j,r}\widetilde{C}_{a-m+j+(q-2N+r)}^{\lambda+N-1-r}$ for a given $r$. If $q-2N+r\leq 0$, then $a-m+j+2(q-2N+r)\leq a-m+j$, so the term has degree $\leq a-m+j$. On the contrary, if q2N+r>0q-2N+r>0, then $D_{j,r}=0$ as $\frac{\Gamma(q-2N+r)}{\Gamma(q-2N)}=0$. Thus, all terms in the sum have degree $\leq a-m+j$. In particular, this expression for $f_{j}$ satisfies (4.2).
@@
In fact, set $\ell:=a+m-2N+j+2r$ and $b:=-\lambda-N+1+r=a+q-2N+r$. Since $q\geq\max(0,N-a+m)$ we have $\ell,b\in{\mathbb{N}}$. Now, if $2b-\ell=a-m+2q-j-2N\geq 0$, then (4.45) holds thanks to (8.10). On the other hand, if $2b-\ell=a-m+2q-j-2N<0$ the right-hand side of (4.45) vanishes. But so does the left-hand side as
@@
By a change of indices, the sum part can be written as
@@
appearing in (4.43) is a well-defined non-zero constant for any $a\geq m-N$, any $j=0,\ldots,N$ and any $\lambda=N+1-a-q\in\Lambda(N,a)$. In fact, if j2N+2q=0j-2N+2q=0, (4.46) equals to $1$. If $j-2N+2q\geq 1$, then (4.46) amounts to
@@
5 The case m<Nm<-N
In this section we recall a duality theorem from [39] that shows the existence of an isomorphism between the F-systems $\text{{Sol}}(\mathfrak{n}_{+};\sigma^{2N+1}_{\lambda},\tau_{m,\nu})$ and $\text{{Sol}}(\mathfrak{n}_{+};\sigma^{2N+1}_{\lambda},\tau_{-m,\nu})$ when $|m|\geq N$. This allows us to deduce the case m<Nm<-N from the case $m>N$ in Theorems 1.3 and 1.5.
@@ 5 The case m<Nm<-N
is a well-defined isomorphism.
With the result above it is clear that we can deduce the case m<Nm<-N from the case $m>N$ in Theorems 1.3 and 1.5. We write the detailed proofs below.
Proof of Theorem 1.3 for m<Nm<-N.
@@
Proof of Theorem 1.5 for m<Nm<-N.
From Proposition 5.1 and the definition of the map $\Phi^{+}_{\text{Sol}}$, the expression of the operator $\mathbb{D}_{\lambda,\nu}^{N,m}$ can be obtained by changing the coordinate $u_{s}$ by the coordinate $(-1)^{s}u_{2N-s}$ (for $s=0,\ldots,2N$), and by switching $i$ by $i$ (or, equivalently, $z$ by $\overline{z}$) in the expression of the operator $\mathbb{D}_{\lambda,\nu}^{N,-m}$. By doing the change of indices given by d=2Ndd=2N-d, we obtain precisely (1.6). ∎
@@ 6 Localness Theorem
In this section we prove the localness theorem (Theorem 1.6). The proof is based on the work [30, 31] and uses several known results about distributions and symmetry breaking operators.
First, recall from [31, Fact 5.1] that there is a natural bijection
@@
$\left(\pi\begin{pmatrix}1&&\\ &g&\\ &&1\end{pmatrix}F_{1}\right)(a)$ $=(\sigma^{2N+1})^{\vee}(g)F_{1}(g^{-1}a),\qquad\text{for }g\in SO(3),$
$\left(\pi(e^{tH_{0}})F_{1}\right)(a)$ $=e^{(\lambda-3)t}F_{1}(e^{-t}a),\qquad\text{for }t\in{\mathbb{R}},$
$\left(\pi(n_{+}(c))F_{1}\right)(a)$ $=F_{1}(a-c),\qquad\text{for }c\in{\mathbb{R}}^{3}.$
@@
One can describe the $P$-action on the second projection space as well. However, in this case, since $U_{2}=N_{-}P/P$ is not $N_{+}$-invariant, the action of $N_{+}$ shall be substituted by its infinitesimal action on the Lie algebra $\mathfrak{n}_{+}$. We omit a detailed description of this action, but it can be verified in [31, Lem. 5.13].
@@
${\mathbb{C}}_{m}(g^{\prime})\circ\mathcal{T}_{1}((g^{\prime})^{-1}y,y_{3})\circ(\sigma^{2N+1})^{-1}(g^{\prime})$ $=\mathcal{T}_{1}(y,y_{3}),$  $\qquad\text{ for all }g^{\prime}\in SO(2),$ (6.3)
$\mathcal{T}_{1}(e^{t}y,e^{t}y_{3})$ $=e^{(\lambda+\nu-3)t}\mathcal{T}_{1}(y,y_{3}),$  $\qquad\text{ for all }t\in{\mathbb{R}},$ (6.4)
$\mathcal{T}_{1}(y-z,y_{3})$ $=\mathcal{T}_{1}(y,y_{3}),$  $\qquad\text{ for all }z\in{\mathbb{R}}^{2}.$ (6.5)
@@
This result can be proved following a similar argument to that of [31, Thm. 3.6].
Recall from [39, Lem. 4.1] that the representation $V^{2N+1}$ can be decomposed as a $SO(2)$-module as follows:
@@
In this section we prove Theorem 7.3, that assures that any non-zero SBO in (1.1) for $|m|>N$ is sporadic (see Definition 7.1 below).
@@
Given $V^{2N+1}\in\widehat{SO(3)}$ and $^{\mathbb{C}}_{m}\in\widehat{SO(2)}$, let $\mathcal{P}\subset{\mathbb{C}}^{2}$ be the set of parameters $(\lambda,\nu)$ such that the space of SBOs (1.1) is non-zero. For a fixed $(N,m)$ and $(\lambda,\nu)\in\mathcal{P}$, we say that an SBO ${\mathbb{T}}_{\lambda,\nu}^{N,m}$ is sporadic with respect to the pair $(SO_{0}(4,1),SO_{0}(3,1))$ if $(\lambda,\nu)$ is an isolated point of $\mathcal{P}$. In other words, if there is an open neighbourhood $U$ of $(\lambda,\nu)$ on ${\mathbb{C}}^{2}$ such that $U\cap\mathcal{P}=\{(\lambda,\nu)\}$.
Although Definition 7.1 above is stated for the pair $(SO_{0}(4,1),SO_{0}(3,1))$ and the representations $V^{2N+1}$ and ${\mathbb{C}}_{m}$, the notion of sporadicity can be formulated also for other pairs $(G,G^{\prime})$ and representations $V,W$ in the obvious way.
@@
    Note that the notion of sporadicity given in [31] by Kobayashi–Speh is slightly different to the one defined above. However, it is straightforward to check that Definition 7.1 implies the one given in [31]. In particular, sporadic SBOs cannot be obtained by residues of meromorphic families of SBOs, contrarily to the regular SBOs (cf. [31]).